# Exact calculation of degrees for lattice equations: a singularity approach

### v2

@@ 4. Proofs
Let $z^{*}\in\mathbb{P}^{1}$ and let $\beta\in\mathbb{P}^{1}(\mathbb{C})$. Then, $z^{*}\in\operatorname{supp}D_{t,n}(\beta)$ if and only if $x_{t,n}(z^{*})=\beta$ and it is the first singularity of the pattern corresponding to $z=z^{*}$. Moreover, the coefficient of $z^{*}$ in $D_{t,n}(\beta)$ coincides with the multiplicity of the constant singularity $x_{t,n}(z^{*})=\beta$.
@@
$\operatorname{div}\left(x_{t,n}=\alpha_{j}\mid\mathbb{P}^{1}\setminus\mathbb{P}^{1}(\mathbb{C})\right)=\sum_{(s,m)\in H}\sum_{i}\operatorname{mult}^{\text{basic},\beta_{i}}_{t-s,n-m}(\alpha_{j})D_{s,m}(\beta_{i}),$
@@
$\mathbb{P}^{1}\setminus\mathbb{P}^{1}(\mathbb{C})=\bigsqcup_{(s,m)\in H}\Delta_{s,m},$
@@
$\operatorname{div}\left(x_{t,n}=\alpha_{j}\mid\mathbb{P}^{1}\setminus\mathbb{P}^{1}(\mathbb{C})\right)$ $=\sum_{(s,m)\in H}\operatorname{div}\left(x_{t,n}=\alpha_{j}\mid\Delta_{s,m}\right)$
@@
If $(s,m)\in H_{0}$, then we have $Z_{s,m}(\beta_{i})=0$ since $Z_{s,m}(\beta_{i})$ corresponds to a movable pattern. Note that for each $(t,n)\in H\setminus H_{0}$, the degree relations at $(t,n)$ consist of $J-1$ linear equations on $Z_{s,m}(\beta_{i})$ ($(s,m)\in H_{\leq(t,n)}\setminus H_{0}$ and $i=1,\ldots,I$). In this proof, we think of $Z_{s,m}(\beta_{i})$ (resp. the degree relations) just as unknown variables (resp. linear equations). We consider whether or not the system of linear equations on these variables has a unique solution.
First, we suppose that $I\geq J$. Let $(t,n)\in H\setminus H_{0}$ and let $r=\#(H_{\leq(t,n)}\setminus H_{0})$. Then, the number of the variables $Z_{s,m}(\beta_{i})$ ($(s,m)\in H_{\leq(t,n)}\setminus H_{0}$ and $i=1,\ldots,I$) is $rI$. On the other hand, the number of the linear equations in this region is $r(J-1)$ since we have $J-1$ equations for each $(s,m)$. Therefore, the system of the linear equations in this region is under-determining and thus the solution is not unique.
@@
  • a movable pattern contains no singular value unless its starting value is one of $\alpha_{1},\ldots,\alpha_{J-1}$,
@@
If $(t_{0},n_{0}+1)\notin H_{0}$, then $(t_{0}-1,n_{0})\in H_{0}$ and
$x_{t_{0},n_{0}+1}(z)=\Phi\left(x_{t_{0}-1,n_{0}+1},z,x_{t_{0}-1,n_{0}}\right).$
Here, we simply used $x_{t_{0}-1,n_{0}+1}$ and $x_{t_{0}-1,n_{0}}$ since $x_{t_{0}-1,n_{0}+1}(z)$ and $x_{t_{0}-1,n_{0}}(z)$ do not depend on $z=x_{t_{0},n_{0}}$. By Proposition 3.17, $x_{t_{0}-1,n_{0}+1}$ depends on the initial variable $x_{t_{0}-t^{\prime},n_{0}+1}$ where $t^{\prime}$ is the least positive integer satisfying $(t_{0}-t^{\prime},n_{0}+1)\in H_{0}$. Since $\Phi(B,x^{*},D)\notin\mathbb{P}^{1}(\mathbb{C})$, $x_{t_{0},n_{0}+1}(x^{*})$ must depend on at least one of the initial variables $x_{t_{0}-1,n_{0}}$ and $x_{t_{0}-t^{\prime},n_{0}+1}$. On the other hand, if $(t_{0},n_{0}+1)\in H_{0}$, then $x_{t_{0},n_{0}+1}$ itself is an initial variable. In both cases, $x_{t_{0},n_{0}+1}(x^{*})$ is not a constant singularity. Interchanging the $t$- and $n$- axes, one can show that $x_{t_{0}+1,n_{0}}(x^{*})$ is not a constant singularity, either.
Since $x_{t_{0},n_{0}+1}(x^{*})$ and $x_{t_{0}+1,n_{0}}(x^{*})$ depend on different initial variables, they are algebraically independent over $\mathbb{C}$. Therefore, using $\Phi(B,C,x^{*})\notin\mathbb{P}^{1}(\mathbb{C})$, we have
$x_{t_{0}+1,n_{0}+1}(x^{*})=\Phi(x_{t_{0},n_{0}+1}(x^{*}),x_{t_{0}+1,n_{0}}(x^{*}),x^{*})\notin\mathbb{P}^{1}(\mathbb{C}).$
That is, $x_{t_{0}+1,n_{0}+1}(x^{*})$ is not a constant singularity. Since none of $x_{t_{0},n_{0}+1}(x^{*})$, $x_{t_{0}+1,n_{0}+1}(x^{*})$ and $x_{t_{0}+1,n_{0}}(x^{*})$ is a constant singularity, it follows from Proposition 3.18 that this pattern has no constant singularity other than $x_{t_{0},n_{0}}(x^{*})=x^{*}$.
The last statement follows from Theorem 3.11 since a basic pattern is a fixed pattern. ∎
@@
Since $\deg_{z}x_{t,n}(z)\geq 1$, there exists $z^{*}\in\mathbb{P}^{1}$ such that $x_{t,n}(z^{*})=\alpha$. Consider the pattern corresponding to $z=z^{*}$. If it is movable, then its starting point is not $(t,n)$ since its starting value is not $\alpha$. If the pattern is fixed, then the starting point is not $(t,n)$ since $(t,n)\in H\setminus H_{0}$. In both cases, the pattern has at least two constant singularities, $x_{t,n}(z^{*})=\alpha$ and the starting point, and is not solitary. ∎
@@
First, suppose that $x_{t,n}(z^{*})=x^{*}$ is the first constant singularity of a movable pattern. We show that the relation $\Phi(B,C,D)=x^{*}$ has a non-trivial factor. By Corollary 4.14, it is impossible for $x_{t,n}(z^{*})$ to become $x^{*}$ as $\frac{0}{0}$. Since none of $x_{t-1,n}(z^{*})$, $x_{t,n-1}(z^{*})$ and $x_{t-1,n-1}(z^{*})$ is $\infty$, there exists a prime factor $F\in\mathbb{C}[B,C,D]$ of the relation
@@
Since $F$ is a prime factor of the relation $\Phi(B,C,D)=x^{*}$, we have
@@ 5. Examples
where $a$ and $b$ are non-zero parameters and $k$ is a positive even integer. This equation was discovered by the present author and reported in detail with collaborators as the first-ever example of a non-integrable lattice equation that passes the singularity confinement test [28, 25]. Let $z=x_{t_{0},n_{0}}$ be an arbitrary initial variable and let
@@
$-B^{k}C^{k}D+aC^{k}+bB^{k}-x^{*}B^{k}C^{k}=0.$
The left-hand side is irreducible since it is degree $1$ with respect to $D$ and the coefficients of $D^{1}$ and $D^{0}$ share no prime factor. Therefore, the equation satisfies the basic pattern condition for $x^{*}$.
@@
it follows from Proposition 3.19 that the basic patterns starting with $x^{\text{basic}}_{t,n}=\beta\notin\{0,\infty\}$ and the fixed patterns corresponding to $z=\gamma\notin\{0,\infty\}$ are solitary. Therefore, it is sufficient to calculate the basic pattern corresponding to $x^{\text{basic}}_{0,0}=0$ and the fixed patterns corresponding to $z=0,\infty$.
@@
where “reg” means that it does not take a constant singularity at $\varepsilon=0$. By Proposition 3.18, this pattern does not have a constant singularity outside the above region. Therefore, by Theorem 3.11, the movable pattern starting with $x_{t,n}(z^{*})=0$ is
@@
[…]
@@
$\ &&\infty^{k}\\ &\infty^{k}&\\ \infty^{k}&&\\ {\color[rgb]{1,0,0}z=0}&{\color[rgb]{1,0,0}x_{t_{0}+1,n_{0}}}\end{matrix}$
@@
$\ &&&\infty^{k}\\ {\color[rgb]{1,0,0}x_{t_{0},n_{0}+1}}&&\infty^{k}\\ {\color[rgb]{1,0,0}z=0}&\infty^{k}\end{matrix},$
@@
$\deg_{z}x_{t,n}(z)$ $=Z_{t,n}(0)+Z_{t-1,n-1}(0)+\delta^{t,1}_{n,0}$
$=kZ_{t-1,n}(0)+kZ_{t,n-1}(0)+\delta_{t-1,n}+k(\delta_{t,n}-\delta^{t,0}_{n,0}),$
where $Z_{t,n}(0)=0$ for $(t,n)\in\mathbb{Z}^{2}\setminus(H\setminus H_{0})$. Solving the linear equation
@@
one finds that the degree growth is exponential in any direction in the first quadrant. If $k=2$, for example, $Z_{t,n}(0)$ and $\deg_{z}x_{t,n}(z)$ are as follows:
@@
and let $z=x_{t_{0},n_{0}}$ be an arbitrary initial variable. Using $\Phi$, we can write the equation as
@@
The singularity structure of this equation has been analyzed in [11, 35], which implies that ±$\pm k$ and ±$\pm\frac{1}{k}$ are important values. However, for a reader who is not familiar with singularity analysis, we start with detecting singularities.
@@
which implies that the equation satisfies the $\partial$-factor condition. Moreover, by Proposition 3.19, one finds that any fixed pattern is solitary unless its starting value is ±$\pm k$ or ±$\pm\frac{1}{k}$, which is why we focus on these four values. The relation $\Phi(B,C,D)=\pm\frac{1}{k}$ can be written as
@@
respectively. Therefore, by Proposition 3.22, the values ±$\pm\frac{1}{k}$ do not appear as the starting value of any movable pattern.
@@
Since it is degree $1$ with respect to $D$, $f$ is irreducible unless $f_{0}$ and $f_{1}$ share a prime factor. Seeking for a contradiction, assume that $f_{1}$ and $f_{2}$ share a prime factor $g$. Since
@@
$g=B-C.$
However, $f_{1}$ is not divisible by $B-C$, which is a contradiction. Therefore, $f$ is irreducible and thus the equation satisfies the basic pattern condition for $x^{*}$.
@@
The left-hand side is irreducible since it is degree $1$ with respect to $D$ and the coefficients of $D^{1}$ and $D^{0}$ do not share a prime factor. Therefore, the equation satisfies the basic pattern condition for $\infty$, too.
We calculate movable patterns starting with ±$\pm k$. A direct calculation shows that the basic patterns corresponding to $x^{\text{basic}}_{0,0}=\pm k$ are
±$\ {\color[rgb]{1,0,0}\bullet}&\pm\frac{1}{k}&\mp k&\circ\\ {\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}x^{\text{basic}}_{0,0}=\pm k}&\mp\frac{1}{k}&\circ\\ &{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}\end{matrix},$
@@
Let us calculate the fixed patterns containing ±$\pm k$ or ±$\pm\frac{1}{k}$. Since any fixed pattern corresponding to $z=\gamma\notin\left\{\pm k,\pm\frac{1}{k}\right\}$ is solitary, it is sufficient to calculate the fixed patterns corresponding to $z=\pm k$, ±$\pm\frac{1}{k}$. A calculation similar to that in Example 5.3 shows that the fixed patterns corresponding to $z=\pm\frac{1}{k}$ do not depend on the shape of $H$:
±$\ &&\pm\frac{1}{k}\\ &\pm\frac{1}{k}\\ {\color[rgb]{1,0,0}z=\pm\frac{1}{k}}\end{matrix}.$
@@
  • ±$\begin{matrix}\pm\frac{1}{k}&\mp k\\ {\color[rgb]{1,0,0}z=\pm k}&\mp\frac{1}{k}\end{matrix}$  (confining),
  • ±$\ &&\pm\frac{1}{k}\\ &\pm\frac{1}{k}\\ \pm\frac{1}{k}\\ {\color[rgb]{1,0,0}z=\pm k}&{\color[rgb]{1,0,0}x_{t_{0}+1,n_{0}}}\end{matrix}$  (non-confining),
  • $\ &&&\mp\frac{1}{k}\\ {\color[rgb]{1,0,0}x_{t_{0},n_{0}+1}}&&\mp\frac{1}{k}\\ {\color[rgb]{1,0,0}z=\pm k}&\mp\frac{1}{k}\end{matrix}$  (non-confining),
@@
where $Z_{t,n}(\pm k)=0$ for $(t,n)\in\mathbb{Z}^{2}\setminus(H\setminus H_{0})$. Each line represents the degree calculated as the number of preimages for $k$, $-k$, $\frac{1}{k}$ and $-\frac{1}{k}$, respectively. It follows from the degree relations that

### v3

@@ 4. Proofs
Let $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ and let $\beta\in\mathbb{P}^{1}(\mathbb{C})$. Then, $z^{*}\in\operatorname{supp}D_{t,n}(\beta)$ if and only if $x_{t,n}(z^{*})=\beta$ and it is the first singularity of the pattern corresponding to $z=z^{*}$. Moreover, the coefficient of $z^{*}$ in $D_{t,n}(\beta)$ coincides with the multiplicity of the constant singularity $x_{t,n}(z^{*})=\beta$.
@@
$\operatorname{div}\left(x_{t,n}=\alpha_{j}\mid\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)\setminus\mathbb{P}^{1}(\mathbb{C})\right)=\sum_{(s,m)\in H}\sum_{i}\operatorname{mult}^{\text{basic},\beta_{i}}_{t-s,n-m}(\alpha_{j})D_{s,m}(\beta_{i}),$
@@
$\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)\setminus\mathbb{P}^{1}(\mathbb{C})=\bigsqcup_{(s,m)\in H}\Delta_{s,m},$
@@
$\operatorname{div}\left(x_{t,n}=\alpha_{j}\mid\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)\setminus\mathbb{P}^{1}(\mathbb{C})\right)$ $=\sum_{(s,m)\in H}\operatorname{div}\left(x_{t,n}=\alpha_{j}\mid\Delta_{s,m}\right)$
@@
If $(s,m)\in H_{0}$, then we have $Z_{s,m}(\beta_{i})=0$ since $Z_{s,m}(\beta_{i})$ corresponds to a movable pattern. Note that for each $(t,n)\in H\setminus H_{0}$, the degree relations at $(t,n)$ consist of J1J-1 linear equations on $Z_{s,m}(\beta_{i})$ ($(s,m)\in H_{\leq(t,n)}\setminus H_{0}$ and $i=1,\ldots,I$). In this proof, we think of $Z_{s,m}(\beta_{i})$ (resp. the degree relations) just as unknown variables (resp. linear equations). We consider whether or not the system of linear equations on these variables has a unique solution.
First, we suppose that $I\geq J$. Let $(t,n)\in H\setminus H_{0}$ and let $r=\#(H_{\leq(t,n)}\setminus H_{0})$. Then, the number of the variables $Z_{s,m}(\beta_{i})$ ($(s,m)\in H_{\leq(t,n)}\setminus H_{0}$ and $i=1,\ldots,I$) is $rI$. On the other hand, the number of the linear equations in this region is r(J1)r(J-1) since we have J1J-1 equations for each $(s,m)$. Therefore, the system of the linear equations in this region is under-determining and thus the solution is not unique.
@@
  • a movable pattern contains none of these singular values unless its starting value is one of $\alpha_{1},\ldots,\alpha_{J-1}$,
@@
Let $x_{t,n}(z^{*})=x^{*}\in\mathbb{C}$ be the first singularity of a pattern.
First, we show that $x_{t,n+1}(z^{*})$ depends on at least one of the variables $x_{t-t^{\prime},n+1}$ and $x_{t-1-t^{\prime\prime},n}$, where $t^{\prime}$ (resp. $t^{\prime\prime}$) is the least nonnegative integer satisfying $(t-t^{\prime},n+1)\in H_{0}$ (resp. $(t-1-t^{\prime},n)\in H_{0}$). If $(t,n+1)\in H_{0}$ (which can occur only if $(t,n)\in H_{0}$), then $x_{t,n+1}$ itself is an initial variable (i.e., $t^{\prime}=0$) and there is nothing to prove. From here on, we consider the case where $(t,n+1)\in H\setminus H_{0}$. By Proposition 3.22, $x_{t-1,n+1}(z^{*})$ depends on $x_{t-t^{\prime},n+1}$. On the other hand, $x_{t-1,n}(z)\in\mathbb{K}_{t-1,n}(z)$ depends on $x_{t-t^{\prime\prime},n}$, and Proposition 4.13 implies that $z^{*}$ is transcendental over $\mathbb{K}_{t-1,n}$. Therefore, $x_{t-1,n}(z^{*})$ depends on $x_{t-t^{\prime\prime},n}$. Since $x_{t-1,n+1}(z^{*})$ and $x_{t-1,n}(z^{*})$ depend on different initial variables, they are algebraically independent and we have
$x_{t,n+1}(z^{*})=\Phi(x_{t-1,n+1}(z^{*}),x^{*},x_{t-1,n}(z^{*})).$
Since it does not belong to $\mathbb{P}^{1}(\mathbb{C})$, $\Phi(B,x^{*},D)$ depends on either $B$ or $D$. Substituting $x_{t-1,n+1}(z^{*})$ and $x_{t-1,n}(z^{*})$ for $B$ and $D$ respectively, we conclude that $x_{t,n+1}(z^{*})$ depends on either $x_{t-t^{\prime},n+1}$ or $x_{t-1-t^{\prime\prime},n}$. Interchanging the $t$- and $n$- axes, one can show that $x_{t+1,n}(z^{*})$ depends on either $x_{t+1,n-n^{\prime}}$ or $x_{t,n-1-n^{\prime\prime}}$, where $n^{\prime}$ (resp. $n^{\prime\prime}$) is the least nonnegative integer satisfying $(t+1,n-n^{\prime})\in H_{0}$ (resp. $(t,n-1-n^{\prime\prime})\in H_{0}$).
Next, we show that $x_{t+1,n+1}(z^{*})$ is not a constant singularity. Since $x_{t,n+1}(z^{*})$ and $x_{t+1,n}(z^{*})$ depend on different initial variables, they are algebraically independent over $\mathbb{C}$. Therefore, using $\Phi(B,C,x^{*})\notin\mathbb{P}^{1}(\mathbb{C})$, we have
$x_{t+1,n+1}(z^{*})=\Phi(x_{t,n+1}(z^{*}),x_{t+1,n}(z^{*}),x^{*})\notin\mathbb{P}^{1}(\mathbb{C}).$
That is, $x_{t+1,n+1}(z^{*})$ is not a constant singularity.
Since none of $x_{t,n+1}(z^{*})$, $x_{t+1,n+1}(z^{*})$ and $x_{t+1,n}(z^{*})$ is a constant singularity, it follows from Proposition 3.23 that this pattern has no constant singularity other than $x_{t,n}(z^{*})=x^{*}$, which completes the proof. ∎
@@
Since $\deg_{z}x_{t,n}(z)\geq 1$, there exists $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ such that $x_{t,n}(z^{*})=\alpha$. Consider the pattern corresponding to $z=z^{*}$. If it is movable, then its starting point is not $(t,n)$ since its starting value is not $\alpha$. If the pattern is fixed, then the starting point is not $(t,n)$ since $(t,n)\in H\setminus H_{0}$. In both cases, the pattern has at least two constant singularities, $x_{t,n}(z^{*})=\alpha$ and the starting point, and is not solitary. ∎
@@
First, suppose that $x_{t,n}(z^{*})=x^{*}$ is the first constant singularity of a movable pattern. We show that the relation $\Phi(B,C,D)=x^{*}$ has a non-trivial factor. By Corollary 4.15, it is impossible for $x_{t,n}(z^{*})$ to become $x^{*}$ as $\frac{0}{0}$. Since none of $x_{t-1,n}(z^{*})$, $x_{t,n-1}(z^{*})$ and $x_{t-1,n-1}(z^{*})$ is $\infty$, there exists an irreducible factor $F\in\mathbb{C}[B,C,D]$ of the relation
@@
Since $F$ is an irreducible factor of the relation $\Phi(B,C,D)=x^{*}$, we have
@@ 5. Examples
Let us consider the polynomial equation
$x_{t,n}=\Phi\left(x_{t-1,n},x_{t,n-1},x_{t-1,n-1}\right)=x_{t,n-1}(x_{t-1,n}-x_{t-1,n-1}).$
Let $X_{t,n}$ be the multiplicity of $x_{t,n}(\infty)=\infty$, i.e.,
$x_{t,n}\left(\varepsilon^{-1}\right)=y_{t,n}\varepsilon^{-X_{t,n}}+O\left(\varepsilon^{-X_{t,n}+1}\right)$
with $y_{t,n}\neq 0$. It follows from the discussion in Example 5.1 that
$\deg_{z}x_{t,n}(z)=X_{t,n}.$
First, we show that $X_{t,n}$ satisfies the ultra-discrete version of the equation:
$X_{t,n}=X_{t,n-1}+\max\left(X_{t-1,n},X_{t-1,n-1}\right).$
The fact that $X_{t,n}$ satisfies this ultra-discrete form is not immediate, because the original equation is not subtraction-free. We now explain why this holds.
If $X_{t-1,n}\neq X_{t-1,n-1}$, then the ultra-discrete equation holds for $(t,n)$ since their leading orders differ. Furthermore, since $x_{t,n}\left(\varepsilon^{-1}\right)$ is a polynomial in $\varepsilon^{-1}$, we have
$X_{t,n}\geq X_{t,n-1}.$
Therefore, using the coefficient $y_{t,n}$ appearing in the above expansion, we have
$x_{t,n}\left(\varepsilon^{-1}\right)=\begin{cases}y_{t,n-1}y_{t-1,n}\varepsilon^{-X_{t,n-1}-X_{t-1,n}}+(\text{lower})&(X_{t,n}>X_{t,n-1}),\\ y_{t,n-1}(y_{t-1,n}-y_{t-1,n-1})\varepsilon^{-X_{t,n-1}-X_{t-1,n}}+(\text{lower})&(X_{t,n}=X_{t,n-1}),\end{cases}$
where $(\text{lower})$” refers to terms of lower order in the expansion with respect to $\varepsilon^{-1}$. It can be shown by induction on $(t,n)$ that
  • $y_{t,n}$ depends on the initial variables $x_{t-t^{\prime},n}$ and $x_{t,n-n^{\prime}}$ where $t^{\prime}$ (resp. $n^{\prime}$) is the least positive integer satisfying $(t-t^{\prime},n)\in H_{0}$ (resp. $(t,n-n^{\prime})\in H_{0}$), i.e., $y_{t,n}\in\Delta_{t,n}$ (see Definition 4.5),
  • $y_{t,n}=\begin{cases}y_{t,n-1}y_{t-1,n}&(X_{t,n}>X_{t,n-1}),\\ y_{t,n-1}(y_{t-1,n}-y_{t-1,n-1})&(X_{t,n}=X_{t,n-1}).\end{cases}$
The key point is that $y_{t-1,n}-y_{t-1,n-1}\neq 0$ since $y_{t-1,n}$ depends on the initial variable $x_{t-t^{\prime},n}$ but $y_{t-1,n-1}$ does not. Therefore, $X_{t,n}$ indeed satisfies the ultra-discrete form. Moreover, because $X_{t-1,n}\geq X_{t-1,n-1}$ for all $(t,n)$, the ultra-discrete equation simplifies to
$X_{t,n}=X_{t,n-1}+X_{t-1,n}.$
Next, we give the solution to the ultra-discrete equation for the following choice of $H$ and $z$:
$H=\mathbb{Z}^{2}_{\geq 0},\quad z=x_{1,0}.$
Since $X_{t,n}=X_{t,n-1}+X_{t-1,n}$, the exact solution is given by Pascal’s triangle and is expressed as
$X_{t,n}=\frac{(t+n-1)!}{(t-1)!\ n!}\quad(t,n\geq 1),$
which yields the following values:
$X_{t,n}=\deg_{z}x_{t,n}(z)\colon\quad\begin{matrix}{{\color[rgb]{1,0,0}0}}&1&7&28&84&210&462&924\\ {{\color[rgb]{1,0,0}0}}&1&6&21&56&126&252&462\\ {{\color[rgb]{1,0,0}0}}&1&5&15&35&70&126&210\\ {{\color[rgb]{1,0,0}0}}&1&4&10&20&35&56&84\\ {{\color[rgb]{1,0,0}0}}&1&3&6&10&15&21&28\\ {{\color[rgb]{1,0,0}0}}&1&2&3&4&5&6&7\\ {{\color[rgb]{1,0,0}0}}&1&1&1&1&1&1&1\\ {{\color[rgb]{1,0,0}0}}&{{\color[rgb]{1,0,0}1}}&{{\color[rgb]{1,0,0}0}}&{{\color[rgb]{1,0,0}0}}&{{\color[rgb]{1,0,0}0}}&{{\color[rgb]{1,0,0}0}}&{{\color[rgb]{1,0,0}0}}&{{\color[rgb]{1,0,0}0}}\end{matrix}.$
The degree exhibits exponential growth, as indicated by the binomial coefficient formula.
Example 5.3.
@@
Example 5.4.
Let us consider the discrete Liouville equation
As discussed in Example 5.3, because the equation has the Laurent property, we can compute the degrees by calculating the fixed patterns corresponding to $z=0,\infty$.
Since the equation is subtraction-free, these patterns can be obtained via ultra-discretization. That is, the leading order of the expansion of $x_{t,n}\left(\varepsilon\right)$ (or $x_{t,n}\left(\varepsilon^{-1}\right)$) with respect to $\varepsilon^{-1}$, denoted by $X_{t,n}$, satisfies the ultra-discrete form
$X_{t,n}=\max\left(X_{t-1,n}+X_{t,n-1},0\right)-X_{t-1,n-1}.$
Let us fix $H$ and $z$ as
$H=\mathbb{Z}^{2}_{\geq 0}\setminus\{(0,0)\},\quad z=x_{1,1}.$
Solving the ultra-discrete equation yields the following patterns:
[…]
and
$\begin{matrix}{\color[rgb]{1,0,0}\vdots}&\vdots&\vdots&\vdots&\iddots\\ {\color[rgb]{1,0,0}x_{0,3}}&\infty&\infty&\infty&\ldots\\ {\color[rgb]{1,0,0}x_{0,2}}&\infty&\infty&\infty&\ldots\\ {\color[rgb]{1,0,0}x_{0,1}}&{\color[rgb]{1,0,0}x_{1,1}=\infty}&\infty&\infty&\ldots\\ &{\color[rgb]{1,0,0}x_{1,0}}&{\color[rgb]{1,0,0}x_{2,0}}&{\color[rgb]{1,0,0}x_{3,0}}&{\color[rgb]{1,0,0}\ldots}\end{matrix}.$
Since the degree coincides with the sum of the multiplicities of $x_{t,n}=\infty$ in these two patterns, we have
$\deg_{z}x_{t,n}(z)=\begin{cases}0&(\min(t,n)=0),\\ 1&(\min(t,n)=1),\\ 2&(\min(t,n)\geq 2).\end{cases}$
This result is in perfect agreement with the degrees previously calculated in [22].
@@
where $a$ and $b$ are non-zero parameters and $k$ is a positive even integer. This equation was discovered by the present author and reported in detail with collaborators as the first-ever example of a non-integrable lattice equation that passes the singularity confinement test [31, 28]. Let $z=x_{t_{0},n_{0}}$ be one of the initial variables and let
@@
$B^{k}C^{k}D+aC^{k}+bB^{k}-x^{*}B^{k}C^{k}=0.$
The left-hand side is irreducible since it is degree $1$ with respect to $D$ and the coefficients of $D^{1}$ and $D^{0}$ share no irreducible factor. Therefore, the equation satisfies the basic pattern condition for $x^{*}$.
@@
it follows from Proposition 3.24 that any pattern whose starting value is neither 0 nor $\infty$ is solitary. Therefore, it is sufficient to calculate the basic pattern corresponding to $x^{\text{basic}}_{0,0}=0$ and the fixed patterns corresponding to $z=0,\infty$.
@@
where “reg” means that it does not have a constant singularity at $\varepsilon=0$. By Proposition 3.23, this pattern does not have a constant singularity outside the above region. Therefore, by Theorem 3.16, the movable pattern starting with $x_{t,n}(z^{*})=0$ is
@@
$\begin{matrix}&&&\iddots\\ &&\infty\\ &\infty\\ {\color[rgb]{1,0,0}z=\infty}\end{matrix}.$
@@
$\begin{matrix}&&&\iddots\\ &&\infty^{k}\\ &\infty^{k}&\\ \infty^{k}&&\\ {\color[rgb]{1,0,0}z=0}&{\color[rgb]{1,0,0}x_{t_{0}+1,n_{0}}}\end{matrix}$
@@
$\begin{matrix}&&&&\iddots\\ &&&\infty^{k}\\ {\color[rgb]{1,0,0}x_{t_{0},n_{0}+1}}&&\infty^{k}\\ {\color[rgb]{1,0,0}z=0}&\infty^{k}\end{matrix},$
@@
$\deg_{z}x_{t,n}(z)$ $=Z_{t,n}(0)+Z_{t-1,n-1}(0)+\delta^{t,1}_{n,0}$ $(\text{$\#$ of preimages of $0$})$
$=kZ_{t-1,n}(0)+kZ_{t,n-1}(0)+\delta_{t-1,n}+k(\delta_{t,n}-\delta^{t,0}_{n,0})$ $(\text{$\#$ of preimages of $\infty$}),$
where $\delta^{a,b}_{c,d}=\delta_{a,b}\delta_{c,d}$ and $Z_{t,n}(0)=0$ for $(t,n)\in\mathbb{Z}^{2}\setminus(H\setminus H_{0})$. From the analysis of the linear equation
@@
one finds that the degree growth is exponential in any direction in the first quadrant. If $k=2$, for example, the following are the results of iterating the recursion over a finite range:
@@
In this paper, we do not address the problem of solving the linear recurrence for $Z_{t,n}(0)$, i.e., finding its exact solution, because this question is not the main focus of the present work and lies within the domain of linear partial difference equations.
@@
and let $z=x_{t_{0},n_{0}}$ be one of the initial variables. Using $\Phi$, we can write the equation as
@@
The singularity structure of this equation has been analyzed in [13, 39], which implies that $\pm k$ and $\pm\frac{1}{k}$ are important values. However, for a reader who is not familiar with singularity analysis, we start with detecting singularities.
@@
which implies that the equation satisfies the $\partial$-factor condition. Moreover, by Proposition 3.24, one finds that any fixed pattern whose starting value is neither $\pm k$ nor $\pm\frac{1}{k}$ is solitary, which is why we focus on these four values. The relation $\Phi(B,C,D)=\pm\frac{1}{k}$ can be written as
@@
respectively. Therefore, by Proposition 3.27, the values $\pm\frac{1}{k}$ do not appear as the starting value of any movable pattern.
@@
Since it is degree $1$ with respect to $D$, $f$ is irreducible unless $f_{0}$ and $f_{1}$ share an irreducible factor. Seeking for a contradiction, assume that $f_{1}$ and $f_{2}$ share an irreducible factor $g$. Since
@@
g=BC.g=B-C.
However, $f_{1}$ is not divisible by BCB-C, which is a contradiction. Therefore, $f$ is irreducible and thus the equation satisfies the basic pattern condition for $x^{*}$.
@@
The left-hand side is irreducible since it is degree $1$ with respect to $D$ and the coefficients of $D^{1}$ and $D^{0}$ do not share an irreducible factor. Therefore, the equation satisfies the basic pattern condition for $\infty$, too.
We calculate movable patterns starting with $\pm k$. A direct calculation shows that the basic patterns corresponding to $x^{\text{basic}}_{0,0}=\pm k$ are
$\begin{matrix}{\color[rgb]{1,0,0}\bullet}&\circ&\circ&\circ\\ {\color[rgb]{1,0,0}\bullet}&\pm\frac{1}{k}&\mp k&\circ\\ {\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}x^{\text{basic}}_{0,0}=\pm k}&\mp\frac{1}{k}&\circ\\ &{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}\end{matrix},$
@@
Let us calculate the fixed patterns containing $\pm k$ or $\pm\frac{1}{k}$. Since any fixed pattern corresponding to $z=\gamma\notin\left\{\pm k,\pm\frac{1}{k}\right\}$ is solitary, it is sufficient to calculate the fixed patterns corresponding to $z=\pm k$, $\pm\frac{1}{k}$. A calculation similar to that in Example 5.5 shows that the fixed patterns corresponding to $z=\pm\frac{1}{k}$ do not depend on the shape of $H$:
$\begin{matrix}&&&\iddots\\ &&\pm\frac{1}{k}\\ &\pm\frac{1}{k}\\ {\color[rgb]{1,0,0}z=\pm\frac{1}{k}}\end{matrix}.$
@@
  • $\begin{matrix}\pm\frac{1}{k}&\mp k\\ {\color[rgb]{1,0,0}z=\pm k}&\mp\frac{1}{k}\end{matrix}$  (confining),
  • $\begin{matrix}&&&\iddots\\ &&\pm\frac{1}{k}\\ &\pm\frac{1}{k}\\ \pm\frac{1}{k}\\ {\color[rgb]{1,0,0}z=\pm k}&{\color[rgb]{1,0,0}x_{t_{0}+1,n_{0}}}\end{matrix}$  (non-confining),
  • $\begin{matrix}&&&&\iddots\\ &&&\mp\frac{1}{k}\\ {\color[rgb]{1,0,0}x_{t_{0},n_{0}+1}}&&\mp\frac{1}{k}\\ {\color[rgb]{1,0,0}z=\pm k}&\mp\frac{1}{k}\end{matrix}$  (non-confining),
@@
where $Z_{t,n}(\pm k)=0$ for $(t,n)\in\mathbb{Z}^{2}\setminus(H\setminus H_{0})$. Each line represents the degree calculated as the number of preimages for $k$, k-k, $\frac{1}{k}$ and $\frac{1}{k}$, respectively. It follows from the degree relations that